# Hardness-Dependent Optimized Quantum Adiabatic Schedules for Maximum Independent Sets

### v1

@@ IV Discussion and outlook
Figure S2: Lower bounds $\Delta_{\text{LB}}$ (purple) and upper bounds $\Delta_{\text{UB}}$ (yellow-green) on the detuning $\Delta_{\text{f}}$ such that the ground state of the Hamiltonian encodes the MIS solution. Calculations have been performed on 10 0001000010\,00010 000 non-isomorphic unit-disk graphs with 8 to 19 vertices. All but 4 graphs satisfy $\Delta_{\text{LB}}\leq V_{0}/12$ (dashed red line) while $\Delta_{\text{UB}}$ is always strictly larger than $V_{0}/8$. It is interesting to note that, in average, easy graphs exhibit much smaller lower bounds than hard graphs (red curve), see also Fig. S1(c-d). Since the average value of $\Delta_{\text{LB}}$ converges quickly with the hardness parameter, we are confident that these bounds can be effectively applied to graphs of any size.
@@
As shown above, the final detuning has to be arbitrarily high for the Hamiltonian to encode in its ground state the MIS solution of some specific graphs. However, this is not the case in general, as most graphs obey well-defined conditions on $\Delta_{\text{f}}$. Based on an extensive analysis of 10 0001000010\,00010 000 graphs, see Fig. S2, we find that more than 99.9% of all graphs satisfy the following bounds:
$\Delta_{\text{f}}\in[\Delta_{\text{LB}},\Delta_{\text{UB}}]=\left[\frac{V_{0}}{12}\,,\frac{V_{0}}{8}\right].$ (S6)
@@
Interestingly, this interval is very close to the best range of parameters found by a variational optimization, see [20, Fig. 6.1], with $\Delta_{\text{LB}}=28.91$ and $\Delta_{\text{UB}}=43.36$ MHz for QuEra’s Aquila device when $a=5$ µm. Note that a statistical analysis is also performed in [18, Sec. S2], where it is stated that about 99% of the MIS problems can be encoded in the ground state of a Rydberg Hamiltonian for well-chosen fixed experimental parameters.
In conclusion, we strongly recommend to consider the interval Eq. (S6) when setting the final detuning of any standard adiabatic protocol aiming to solve the MIS problem. A value which lies far outside this interval will likely lead to poor numerical or experimental results simply because the ground state of the Rydberg Hamiltonian does not correspond to the MIS solution.
SII Graph-dependent counterdiabatic driving
@@
$H(t)=\frac{\Omega(t)}{2}\sum_{i}\hat{X}_{i}-\Delta(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}V_{ij}\hat{n}_{i}\hat{n}_{j}\equiv\frac{\Omega(t)}{2}\ \hat{X}-\Delta(t)\ \hat{n}+\hat{V},$ (S7)
@@
SII.1 Approximate adiabatic gauge potential
In order to suppress the transitions between eigenstates during the shorten adiabatic evolution, the gauge potential $\mathcal{A}$ in Eq. (5) must satisfy the following equation [26, Eq. (2)]:
$\left[\mathrm{i}\mkern 1.0mu\partial_{t}H-\left[\mathcal{A},H\right],H\right]=0.$ (S8)
Neglecting the interactions $\hat{V}$, it is shown in [27] that an exact solution of this equation exists:
$\mathcal{A}_{V\rightarrow 0}=-\frac{1}{2}\ \frac{\Omega\dot{\Delta}-\dot{\Omega}\Delta}{\Omega^{2}+\Delta^{2}}\ \hat{Y}.$ (S9)
This gauge potential leads to improved results compared to a simple adiabatic protocol, but one might expect even better results with a gauge potential that explicitly depends on the coefficients $V_{ij}$, hence being specific to the graph to be solved. In order to find such a generalized gauge potential, we consider the variational procedure that consists in minimizing the action [26, Eq. (8)]
$\mathcal{S}(\mathcal{A}_{\lambda})=\operatorname{Tr}\left(G^{2}(\mathcal{A}_{\lambda})\right)\quad\text{with}\quad G(\mathcal{A}_{\lambda})=\partial_{t}H+\mathrm{i}\mkern 1.0mu\ [\mathcal{A}_{\lambda},H].$ (S10)
@@
$\lambda=2\,\frac{\operatorname{Tr}\left(\partial_{t}H\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},H]\right)}{\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},H]^{2}\right)}.$ (S11)
Based on the algebraic properties of the Pauli matrices and of the trace operator, this expression can be developed and simplified as follows:
$\lambda=\frac{\Omega\dot{\Delta}-\dot{\Omega}\Delta+\dot{\Omega}\ T_{1}}{\Omega^{2}+\Delta^{2}-2\Delta\ T_{1}+T_{2}},$ (S12)
@@
$N\cdot T_{1}$ $=2^{-N}\operatorname{Tr}\left(\hat{X}\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]\right)=\sum_{i<j}V_{ij},$ (S13a)
$N\cdot T_{2}$ $=2^{-N}\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]^{2}\right)=\frac{1}{2}\sum_{i\leq j}\sum_{k\neq i,j}V_{ik}V_{kj}.$ (S13b)
As expected, the graph-dependent gauge potential $\mathcal{A}_{\lambda}$ is equal to $\mathcal{A}_{V\rightarrow 0}$ in the limiting case of vanishing interactions, with $T_{1}=T_{2}=0$. Due to the fast-decaying interaction between distant atoms, $T_{1}$ and $T_{2}$ do not scale with $N$ as they represent average values of (nearly) local functions. For instance, $T_{1}$ is exactly half the mean degree of the underlying graph if ones considers its adjacency matrix instead of the interaction terms in Eq. (S13a).
Finally, the driving amplitude and phase of the CD Hamiltonian as defined in Eq. (5) become:
$\Omega_{\text{CD}}=\sqrt{\Omega^{2}+\lambda^{2}},$ (S14a)
$\phi_{\text{CD}}=\arctan(\lambda/\Omega).$ (S14b)
Note that $\lambda$ and therefore $\dot{\Omega}$ should vanish at the beginning and at the end of the protocol to satisfy the boundary conditions $\Omega_{\text{CD}}(t_{\text{i}})=\Omega_{\text{CD}}(t_{\text{f}})=0$.
SII.2 Practical application of the counterdiabatic driving
@@
It follows from Eq. (S14a) that $\Omega_{\text{CD}}$ is always larger than $\Omega$, so that it may exceed the maximum Rabi frequency $\Omega_{\max}$ of the quantum computer. In this case, we rescale the adiabatic function $\Omega$ by a factor $10<k<$ 1 to satisfy the hardware constraint at any time $t$ of the evolution:
$\max_{t}\left\{\Omega_{\text{CD}}(k\,\Omega(t))\right\}=\Omega_{\max}.$
Note that this is slightly different from rescaling $\Omega_{\text{CD}}$ directly, as the relations (S12, S14) would not hold in that case.
@@
Since Eq. (S12) is a generalization of the counterdiabatic term proposed in [27] through the addition of $T_{1}$ and $T_{2}$, it is interesting in an optimization context to consider a continuous interpolation between these two protocols. One possibility is to consider the interaction strength $V_{0}$ as a free parameter in the range $[0,C_{6}/a^{6}]$ when calculating the CD driving. To avoid any confusion with the fixed interactions in the main Hamiltonian, we denote by $\nu$ this free parameter. Numerically, this is equivalent to replace $T_{1}$ and $T_{2}$ in Eq. (S12) by the following expressions:
$T_{1}$ $=\nu\,T_{1}^{(0)},$ (S15a)
$T_{2}$ $=\nu^{2}\,T_{2}^{(0)},$ (S15b)
where $T_{1}^{(0)}$ and $T_{2}^{(0)}$ are the traces evaluated at $V_{0}=1$ in Eq. (S13).
SIII Graph datasets and numerical implementation
In the following sections, we describe two sets of unit-disk graphs $G=(V,E)$ that are used in this study; data are available on GitHub [41]. In all examples, we consider an underlying square lattice where the distance between nearest-neighbors is $a=1$ and where any pair of vertices in $V$ is connected by an edge in $E$ if the corresponding distance is smaller than or equal to $\sqrt{2}$. Then, we detail the numerical method that is used to simulate the evolution of the quantum state under the action of the Rydberg Hamiltonian. Finally, we provide the optimized schedule functions for the linear and for the counterdiabatic protocol with 6 and 3 free variables, respectively.
SIII.1 Graph datasets
@@
From 10 0001000010\,00010 000 non-isomorphic graphs randomly generated, we choose 500 graphs such that:
@@
The result is shown in Fig. S3. The main idea behind these rules is to generate a set of graphs for which the correlation between the hardness parameter and the order is as small as possible. In fact, in this study we want to show that the optimum parameters for the adiabatic evolution depend strongly on the hardness of the graphs and only weakly on their order.
Figure S3: Visualization of the hardness parameters $\mathcal{HP}$ and the orders of the 500 small graphs. Due to the combinatorial definition of the hardness parameter, graphs with fewer vertices (down purple triangles) cannot generate very hard MIS instances. This explains why graphs of order 13 and more (up yellow-green triangles) are overrepresented in the range $\mathcal{HP}\in[6,12]$.
@@
For completeness, we display in Fig. S4 the 11 “toy graphs” as defined in [28, Appendix C]:
@@
  • there is only one MIS solution
Figure S4: The 11 “toy graphs” sorted by the number of vertices (9 or 10) and by hardness parameter (between 2 and 3). The unique MIS is highlighted with dark blue circles. To refer to a specific graph, we generate the three-nibble identifiers as follows: the $i$-th nibble (four-bit aggregation) is the hexadecimal symbol that represents the absence or presence of a vertex in the 4 sites of the $i$-th row (from top to bottom).
SIII.2 Numerical implementation
Several software development kits (SDK) exist to simulate the Rydberg evolution: Bloqade [44] and Amazon Braket [31] for QuEra’s hardware, or Pulser [32] based on QuTiP for Pasqal’s technology, just to name a few. Nevertheless, we developed our own routines to get a full control over all parameters and, more importantly, to fit exactly our needs. In fact, optimizing hundreds of graphs that consist of $8\leq N\leq 17$ vertices for several protocols requires intense (classical) computing power, so that we have to find the perfect balance between accuracy and speed. For example, for each graph of the datasets we precompute all time- and protocol-independent terms in the Hamiltonian, namely the $\hat{X}$, $\hat{Y}$, $\hat{n}$ and $\hat{V}$ operators, see Eq. (S7). We also use an explicit Runge-Kutta method of order 2 [45] to integrate the time-dependent Schrödinger equation, which turns out to be sufficiently accurate and much faster than the default order 8 method (DOP853 in Python or Vern8 in Julia).
Figure S5: Numerical simulations of the Rydberg Hamiltonian for the toy graph “7DE” and the ACQC protocol defined in [27] with $t_{\text{f}}=1$ µs. The height of the bars represents the probability to find the atom $i$ in the state |$\ket{1}$ at the end of the evolution, with $i=1,\ldots,9$ from top to bottom and left to right in the graph. The MIS solution (1, 3, 7, 9) is clearly visible in this bar chart. We verify that our implementation (light colors) is perfectly equivalent to that of Bloqade (dark colors); note that, depending on the SDK, the phase has to be multiplied by minus one in Eq. (S14b). We also observe that the simulations in the full Hilbert space (blue bars) match those performed in the blockade subspace of nearest neighbors (orange bars). However, the unit-disk subspace approximation $R_{\text{s}}=R_{\text{b}}$ (green bars) is clearly in favor of the MIS solution and thus cannot be considered as satisfactory. This is even more noticeable when comparing the probability to reach the ground state at the end of the evolution for each blockade subspace: 0.49(4), 0.49(6), and 0.70(0) respectively.
@@
Besides the above-mentioned implementation aspects, a crucial ingredient is the correct use of the blockade subspace: when the interaction energy between Rydberg states is much larger than the Rabi strength, the exact dynamics in the full Hilbert space is well approximated by the dynamics in the subspace where only one Rydberg excitation is allowed between nearby atoms [46]. In numerical simulations, the states |$\ket{11}_{ij}$ are therefore neglected if the distance between $i$ and $j$ is smaller than a given subspace radius $R_{\text{s}}$, which results in much faster computations. The default setting in most programs is $R_{\text{s}}=R_{\text{b}}$, so that the quantum evolution is restricted to the independent sets in unit-disk graphs. However, this approximation is too coarse to get accurate results, see Fig. S5: neglecting the states |$\ket{11}$ for next-nearest neighbors (extremities of a diagonal) leads to success probabilities that are overestimated by up to 50%. Therefore, it is crucial to simulate either the full Hilbert space or the blockade subspace where only nearest neighbor Rydberg states are neglected.
@@
In this study, we focus on the probability to reach a MIS at the end of the adiabatic evolution and not on the normalized mean energy (approximation ratio) of the final state, as it often the case for this problem. In fact, all results here revolve around the hardness parameter, which is intrinsically related to the two lowest-energy levels and not to the full energy spectrum. We therefore define the score function to be maximized as the overlap between the quantum state |$\ket{\psi}$ at the end of the evolution and all MIS solutions of the graph under investigation:
$P_{\text{MIS}}(\Theta)=\sum_{i=1}^{D_{|\text{MIS}|}}\big{|}\,\braket{\psi(\Theta,t_{\text{f}})\,}{\,e_{i}}\,\big{|}^{2},$ (S16)
where $e_{i}$ are the vectors of the canonical basis that correspond to the $D_{|\text{MIS}|}$ configurations of the (possibly) degenerate solution. The parameters $\Theta$ to be optimized depend on the protocol, e.g., $\Theta=(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for the simple linear schedules. Since the evolution is deterministic, any standard optimization algorithm may be applied. We choose the simplex search method [47], which turns out to be robust and very efficient in our setting. We set $t_{\text{f}}=1$ µs, which leads to success probabilities that are well distributed over the whole unit interval for the 500 graphs of the dataset.
SIII.3 Optimized adiabatic schedules
The optimum parameters $\Theta$ depend principally on the hardness parameter of the graphs—and only marginally on their order $N$ for instance—so that an effective model $\hat{\Theta}(\mathcal{HP})$ can be fitted to the 500 small graphs. The simplest linear schedules are illustrated in Fig. 2 in the main text. In Fig. S6 and Fig. S7, we display the optimized schedules for a more advanced linear protocol with 6 free parameters and for the generalized counterdiabatic protocol with 3 free parameters. Finally, we benchmark these hardness-dependent protocols against those described in [27]: the success probability is increased for each of the 500 graphs of the dataset, see Fig. S8.
Figure S6: Optimized piecewise linear schedules. The schedules depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon) and in the limit $\mathcal{HP}\to\infty$ (dark maroon thick line). In this latter case, the optimized parameters are $\tau_{\text{i}}\approx 0.22\,t_{\text{f}}$ and $\tau_{\text{f}}\approx 0.12\,t_{\text{f}}$ for the Rabi amplitude (left) and $\tau_{\text{m}}\approx 0.34\,t_{\text{f}}$, $\Delta_{\text{i}}\approx-0.83\,\Delta_{\text{UB}}$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$ for the detuning (right).
Figure S7: Optimized generalized counterdiabatic schedules. In this case, the functions $\Omega$ and $\phi$ depend not only on the hardness parameter but also on the duration of the protocol (here, $t_{\text{f}}=1$ µs) and on the value of the traces $T_{1}$ and $T_{2}$, see Eq. (S12) and Eq. (S14). The thin lines are based on the average values of the 500 small graphs: $\overline{T}_{1}\approx 0.93$ and $\overline{T}_{2}\approx 1.46$. The light green area represents the maximum variation of the phase for a fixed hardness parameter $\mathcal{HP}=0.5$.
Figure S8: Probability of finding a MIS in the 500 small graphs for the main protocols. Compared to the benchmark protocols, $P_{\text{MIS}}$ is improved in average by a factor of 3.6 for the piecewise linear schedules (left) and by a factor of 1.5 for the couterdiabatic schedules (center) developed in this study. Except for rare instances, the $\mathcal{HP}$-linear protocol always beats the $\mathcal{HP}$-ACQC protocol (right). Note that $P_{\text{MIS}}$ depends strongly on $\mathcal{HP}$ but not on $N$, with low values for the hard graphs (maroon) and nearly perfect results for the easy ones (green).

### v2

@@ A.2 Statistical analysis of the bounds on the detuning
Figure 11: Box plot of the lower bounds $\Delta_{\text{LB}}$ (purple) and upper bounds $\Delta_{\text{UB}}$ (yellow-green) on the detuning $\Delta_{\text{f}}$ such that the ground state of the Hamiltonian encodes the MIS solution. Calculations have been performed on 10 0001000010\,00010 000 non-isomorphic unit disk graphs with 8 to 19 vertices. All but 4 graphs satisfy $\Delta_{\text{LB}}\leq V_{0}/12$ (dashed red line) while $\Delta_{\text{UB}}$ is always strictly larger than $V_{0}/8$. It is interesting to note that, in average, easy graphs exhibit much smaller lower bounds than hard graphs (red curve), see also Fig. 10(c-d). Since the average value of $\Delta_{\text{LB}}$ quickly saturates for not-too-easy graphs, we are confident that these bounds can be effectively applied to graphs of any size.
@@
As shown above, the final detuning has to be arbitrarily high if one wants to encode the MIS solution of any graph in the Rydberg Hamiltonian. However, this is not the case in general, as most graphs obey well-defined conditions on $\Delta_{\text{f}}$. Based on an extensive analysis of 10 0001000010\,00010 000 graphs, see Fig. 11, we find that more than 99.9% of all graphs satisfy the following bounds:
$\Delta_{\text{f}}\in[\Delta_{\text{LB}},\Delta_{\text{UB}}]=\left[\frac{V_{0}}{12}\,,\frac{V_{0}}{8}\right].$ (14)
@@
Interestingly, this interval is very close to the best range of parameters found by a variational optimization, see [21, Fig. 6.1], with $\Delta_{\text{LB}}=28.91$ and $\Delta_{\text{UB}}=43.36$ MHz for QuEra’s Aquila device and $a=5$ µm. Note that a statistical analysis is also performed in [19, Sec. S2], where it is stated that about 99% of the MIS problems can be encoded in the ground state of a Rydberg Hamiltonian for well-chosen fixed experimental parameters. We strongly recommend to consider the interval in Eq. (14) when setting the final detuning of any adiabatic protocol for the MIS problem. A value which lies far outside this interval will likely lead to poor numerical or experimental results simply because the ground state of the Hamiltonian does not correspond to the MIS solution.
Appendix B Graph-dependent counterdiabatic driving
@@
$H(t)=\frac{\Omega(t)}{2}\sum_{i}\hat{X}_{i}-\Delta(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}V_{ij}\hat{n}_{i}\hat{n}_{j}\equiv\frac{\Omega(t)}{2}\ \hat{X}-\Delta(t)\ \hat{n}+\hat{V},$ (15)
@@
B.1 Approximate adiabatic gauge potential
The idea of the CD driving is to add a gauge potential $\mathcal{A}$ to the Hamiltonian $H$ such that the transitions between eigenstates are suppressed during a shorten adiabatic evolution:
$\widetilde{H}(t)=H(t)+\mathcal{A}(t),$ (16)
where $\mathcal{A}$ satisfies the following equation [27, Eq. (2)]:
$\left[\mathrm{i}\mkern 1.0mu\partial_{t}H-\left[\mathcal{A},H\right],H\right]=0.$ (17)
Neglecting the interactions $\hat{V}$, it is shown in [28] that a solution of this equation exists:
$\mathcal{A}_{V\rightarrow 0}=-\frac{1}{2}\ \frac{\Omega\dot{\Delta}-\dot{\Omega}\Delta}{\Omega^{2}+\Delta^{2}}\ \hat{Y}.$ (18)
This gauge potential leads to improved results compared to a simple adiabatic protocol, but one might expect even better results with a gauge potential that explicitly depends on the coefficients $V_{ij}$, hence being specific to the graph to be solved. In order to find such a generalized gauge potential, we consider the variational procedure that consists in minimizing the action [27, Eq. (8)]
$\mathcal{S}(\mathcal{A}_{\lambda})=\operatorname{Tr}\left(G^{2}(\mathcal{A}_{\lambda})\right)\quad\text{with}\quad G(\mathcal{A}_{\lambda})=\partial_{t}H+\mathrm{i}\mkern 1.0mu\ [\mathcal{A}_{\lambda},H].$ (19)
@@ B.1 Approximate adiabatic gauge potential
$\lambda=2\,\frac{\operatorname{Tr}\left(\partial_{t}H\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},H]\right)}{\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},H]^{2}\right)}.$ (20)
Based on the algebraic properties of the Pauli matrices and the trace operator, this expression can be developed and simplified as follows:
$\lambda=\frac{\Omega\dot{\Delta}-\dot{\Omega}\Delta+\dot{\Omega}\ T_{1}}{\Omega^{2}+\Delta^{2}-2\Delta\ T_{1}+T_{2}},$ (21)
@@
$N\cdot T_{1}$ $=2^{-N}\operatorname{Tr}\left(\hat{X}\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]\right)=\sum_{i<j}V_{ij},$ (22a)
$N\cdot T_{2}$ $=2^{-N}\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]^{2}\right)=\frac{1}{2}\sum_{i\leq j}\sum_{k\neq i,j}V_{ik}V_{kj}.$ (22b)
As expected, the graph-dependent gauge potential $\mathcal{A}_{\lambda}$ is equal to $\mathcal{A}_{V\rightarrow 0}$ in the case of vanishing interactions, with $T_{1}=T_{2}=0$ in this limit. Due to the fast-decaying interaction between distant atoms, $T_{1}$ and $T_{2}$ do not scale with $N$ as they represent average values of (nearly) local functions. For instance, $T_{1}$ is exactly half the mean degree of the underlying graph if ones considers its adjacency matrix instead of the interaction terms in Eq. (22a).
Finally, the driving amplitude and phase of the full Rydberg Hamiltonian $\widetilde{H}$ become:
$\widetilde{\Omega}=\sqrt{\Omega^{2}+\lambda^{2}},$ (23a)
$\widetilde{\phi}=\arctan(\lambda/\Omega).$ (23b)
Note that $\lambda$ and therefore $\dot{\Omega}$ should be equal to zero at the beginning and at the end of the protocol to satisfy the boundary conditions $\widetilde{\Omega}(t_{\text{i}})=\widetilde{\Omega}(t_{\text{f}})=0$.
B.2 Practical application of the counterdiabatic driving
@@
It follows from Eq. (23a) that $\widetilde{\Omega}$ is always larger than $\Omega$, so that it may exceed the maximum Rabi frequency $\Omega_{\max}$ of the quantum computer. In this case, we rescale the adiabatic function $\Omega$ by a factor $10<\kappa<$ 1 to satisfy the hardware constraint at any time $t$ of the evolution:
$\max_{t}\left\{\widetilde{\Omega}\big{(}\kappa\,\Omega(t),\Delta,T_{1},T_{2}\big{)}\right\}=\Omega_{\max}.$
Note that this is slightly different from rescaling $\widetilde{\Omega}$ directly, as the relations (21, 23) would not hold in that case.
@@
Since Eq. (21) is a generalization of the counterdiabatic term proposed in [28] through the addition of $T_{1}$ and $T_{2}$, it is interesting in an optimization context to consider a continuous interpolation between these two protocols. One possibility is to consider the interaction strength $V_{0}$ as a free parameter in the range $[0,C_{6}/a^{6}]$ when calculating the CD driving. To avoid any confusion with the fixed interactions in the Rydberg Hamiltonian, we denote by $\nu$ this free parameter. Numerically, this is equivalent to replace $T_{1}$ and $T_{2}$ in Eq. (21) by the following expressions:
$T_{1}$ $=\nu\,T_{1}^{(0)},$ (24a)
$T_{2}$ $=\nu^{2}\,T_{2}^{(0)},$ (24b)
where $T_{1}^{(0)}$ and $T_{2}^{(0)}$ are the traces evaluated at $V_{0}=1$ in Eq. (22). Another possibility would be, e.g., to consider the two-variable parametrization $T_{i}=\nu_{i}\,T_{i}^{(0)}$, with $\nu_{i}\in[0,V_{0}^{i}]$ and $i\in\{1,2\}$.
Appendix C Graph datasets and numerical implementation
In the following sections, we first describe two sets of unit disk graphs $\mathcal{G}=(V,E)$ that are used in this study; the data are available on GitHub [40]. In all examples, we consider an underlying square lattice where the distance between nearest-neighbors is $a=1$ and where any pair of vertices in $V$ is connected by an edge in $E$ if the corresponding distance is smaller than or equal to $\sqrt{2}$. Then, we detail the numerical method that is used to simulate the evolution of the quantum state under the action of the Rydberg Hamiltonian.
C.1 Graph datasets
@@
From 10 0001000010\,00010 000 non-isomorphic graphs randomly generated, we choose 500 graphs such that the correlation between their hardness parameter and their order is as small as possible, see Fig. 12. More precisely, we apply the following rules to determine which graphs belong to the dataset:
@@
Figure 12: Visualization of the hardness parameters $\mathcal{HP}$ and the orders of the 500 small graphs. Due to the combinatorial definition of the hardness parameter, graphs with fewer vertices (down purple triangles) cannot generate very hard MIS instances. This explains why graphs of order 13 and more (up yellow-green triangles) are overrepresented in the range $\mathcal{HP}\in[6,12]$.
@@
For completeness, we display in Fig. 13 the 11 “toy graphs” defined in [29, Appendix C]:
@@
  • there is only one MIS solution for each graph
Figure 13: The 11 “toy graphs” sorted by the number of vertices (9 or 10) and by hardness parameter (between 2 and 3). The unique MIS is highlighted with dark blue circles. To refer to a specific graph, we generate the three-nibble identifiers as follows: the $i$-th nibble (four-bit aggregation) is the hexadecimal symbol that represents the absence or presence of a vertex in the 4 sites of the $i$-th row (from top to bottom).
C.2 Numerical implementation
Several software development kits (SDK) exist to simulate the Rydberg evolution: Bloqade [43] and Amazon Braket [32] for QuEra’s hardware, or Pulser [33] based on QuTiP for Pasqal’s technology, just to name a few. Nevertheless, we developed our own routines to get a full control over all parameters and, more importantly, to fit exactly our needs. In fact, optimizing hundreds of graphs for several protocols requires intense (classical) computing power, so that we have to find the perfect balance between accuracy and speed. For example, for each graph of the datasets we precompute all time- and protocol-independent terms in the Hamiltonian, namely the $\hat{X}$, $\hat{Y}$, $\hat{n}$ and $\hat{V}$ operators, see Eq. (15). We also use an explicit Runge-Kutta method of order 2 [44] to integrate the time-dependent Schrödinger equation, which turns out to be sufficiently accurate and much faster than the default order 8 method (DOP853 in Python or Vern8 in Julia). The initial state of this ordinary differential equation is |$\ket{0}^{\otimes N}$, which corresponds to the unique ground state of the Rydberg Hamiltonian with $\Omega=0$ and $\Delta<0$.
@@ C.2 Numerical implementation
Besides the above-mentioned implementation aspects, a crucial ingredient is the correct use of the blockade subspace: when the interaction energy between Rydberg states is much larger than the Rabi strength, the exact dynamics in the full Hilbert space is well approximated by the dynamics in the subspace where only one Rydberg excitation is allowed between nearby atoms [45]. In numerical simulations, the states |$\ket{11}_{ij}$ are therefore neglected if the distance between $i$ and $j$ is smaller than a given subspace radius $R_{\text{s}}$, which results in much faster computations. The default setting in most programs is $R_{\text{s}}=R_{\text{b}}$, so that the quantum evolution is restricted to the independent sets in unit disk graphs. However, this approximation is too coarse to get accurate results, see Fig. 14: neglecting the states |$\ket{11}$ for next-nearest neighbors (extremities of a diagonal) leads to success probabilities that are overestimated by up to 50%. Therefore, it is crucial to simulate either the full Hilbert space or the blockade subspace where only nearest neighbor Rydberg states are neglected.
@@
In this study, we focus on the probability to reach a MIS, i.e., we want to maximize the overlap between the quantum state at the end of the evolution and all MIS solutions of the graph under investigation, see Eq. (5). The set of parameters $\Theta$ to be optimized depends on the protocol, e.g., $\Theta=(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for the simple linear schedules. Since the Schrödinger evolution is deterministic and that we simulate noiseless systems, any standard optimization algorithm may be applied. We choose the simplex search method [46], which turns out to be robust and very efficient in our setting. Finally, we set $t_{\text{f}}=1$ µs, which leads to success probabilities that are well distributed over the whole unit interval for the 500 graphs of the dataset.
Figure 14: Numerical simulations of the Rydberg Hamiltonian for the toy graph “7DE” and the ACQC protocol defined in [28] with $t_{\text{f}}=1$ µs. The height of the bars represents the probability to find the atom $i$ in the state |$\ket{1}$ at the end of the evolution, with $i=1,\ldots,9$ from top to bottom and left to right in the graph. The MIS solution (1, 3, 7, 9) is clearly visible in this bar chart. We verify that our implementation (light colors) is perfectly equivalent to that of Bloqade (dark colors); note that, depending on the SDK, the phase has to be multiplied by minus one in Eq. (23b). We also observe that the simulations in the full Hilbert space (blue bars) match those performed in the blockade subspace of nearest neighbors (orange bars). However, the unit disk subspace approximation $R_{\text{s}}=R_{\text{b}}$ (green bars) is clearly in favor of the MIS solution and thus cannot be considered as satisfactory. This is even more noticeable when comparing the probability to reach the ground state at the end of the evolution for each blockade subspace: 0.49(4), 0.49(6), and 0.70(0) respectively.